OpenStudy (anonymous):

Find the surface area that results when the curve is rotated around the x-axis: x = t ^(2) + 1 , y = t ; 0 ≤ t ≤ 2 .

OpenStudy (anonymous):

????

myininaya (myininaya):

use int(2*pi*y*sqrt((x')^2+(y')^2),t)

myininaya (myininaya):

int(2*pi*t*sqrt(4t^2+1),t)

myininaya (myininaya):

let u=4t^2+1 so du=8tdt so we have int(2/8*u^(1/2),u) that should help you

OpenStudy (anonymous):

ok i'm tryin it now

OpenStudy (amistre64):

would this be equivalent to: x = y^2 +1 ?

OpenStudy (anonymous):

im still workin on it ha

OpenStudy (amistre64):

i spose the line integral dont care if its parametric or not eh :)

OpenStudy (anonymous):

what im confused ha

OpenStudy (amistre64):

the line integral;\[\int\limits_{} \sqrt{f'(x) + g'(x)} dx\]

OpenStudy (anonymous):

i dont know what exactly you're asking but we're doing parametric stuff now?

OpenStudy (anonymous):

haha i'm so confused with this stuff right now sorry

OpenStudy (amistre64):

the line integral is primed for parametric equations; you insert the derivatives of your parameter equations and run with it :)

OpenStudy (anonymous):

so i should find the derivative of both and plug them into the integral you put earlier instead of what i just did?

OpenStudy (amistre64):

dunno what you just did :) but yeah. you find the length of the line, and spin it to find the surface area...

OpenStudy (amistre64):

\[2\pi \int\limits_{0}^{2}\sqrt{f'(t) + g'(t)} dt\]

OpenStudy (amistre64):

if im wrong, im sure someone will let me know :)

OpenStudy (amistre64):

x = t ^(2) + 1 , y = t x' = 2t ; y' = 1 fill em in

OpenStudy (amistre64):

and I was wrong...those whoud be "squared" lol

OpenStudy (amistre64):

\[2\pi \int\limits_{0}^{2}\sqrt{[f'(x)]^{2} + [g'(x)]^2}\]

OpenStudy (anonymous):

ok so i ended up with pi/4(2u^(3/2)/3) with the limits from 1 to 17 bc i plugged in 0 and 2 into u = 4t^2 +1

OpenStudy (anonymous):

and just going to plug in those limits to find out the answer is that right?

OpenStudy (amistre64):

[S] (4t^2 + 1)^(1/2) dt how did you go about integrating the square root? did you use a trig substitutiion? or a "u" substitution?

OpenStudy (anonymous):

u=4t^2 +1

OpenStudy (anonymous):

so du = 8tdt

OpenStudy (amistre64):

and how does that integrate?...good, and how did you get rid of the "t" from that?

OpenStudy (anonymous):

ohhhhh i guess i messed that up

OpenStudy (anonymous):

ahh now im confused again

OpenStudy (amistre64):

\[\int\limits_{}8t (\sqrt{u}) du =?\]

OpenStudy (amistre64):

itd be in the denominator, but still....its there lol

OpenStudy (amistre64):

trig identities are what you can use for substitution in this

OpenStudy (amistre64):

\[\sqrt{4x^2 + 1}\] tan^2 + 1 = sec^2 right?

OpenStudy (amistre64):

say u = 2x tan(x) u^2 = 4x^2 tan^2....plays a role in this someplace :)

OpenStudy (anonymous):

im not gettin it ha

OpenStudy (amistre64):

you pretty much want to make it looke like this:\[\sqrt{\tan^2(x) + 1} = \sec(x)\]

OpenStudy (amistre64):

so that tan(x) = 2x

OpenStudy (amistre64):

then you integrate sec(x) dx and turn the substitution around

OpenStudy (anonymous):

so would it be (2tan^2(x) +1)^1/2

OpenStudy (anonymous):

when i start the trig sub or what?

OpenStudy (amistre64):

no, then you cant "factor" out that "2" in front of the tan... just my thoughts :)

OpenStudy (anonymous):

ahhh ive been workin at this too long i cant think straight now im forgetting how to do trig sub stuff

OpenStudy (amistre64):

4x^2 needs to equal tan^2 2x = tan then

OpenStudy (amistre64):

we draw a triangle with the corresponding parts so that one leg = 2x; the other leg = 1 and the hypotenuse = whatever it equals :)

OpenStudy (amistre64):

that way we can convert between our stuff in the end right?

OpenStudy (anonymous):

ya i mean if we did the triangle it would be sqrt of 2x^2 + 1^1

OpenStudy (anonymous):

ok so heres where im at

OpenStudy (amistre64):

does this help?

OpenStudy (anonymous):

2 pi (ln(sec(t) + tan(t)) and whatever the limits are

OpenStudy (anonymous):

does that seem correct?

OpenStudy (amistre64):

that looks good, just need to replace sec and tan with their appropriate values

OpenStudy (anonymous):

ok so tan is 2x right?

OpenStudy (amistre64):

correct.... or 2t...whatever you wanna call it :)

OpenStudy (anonymous):

and sec is adj/opp? so 1/2x

OpenStudy (anonymous):

?

OpenStudy (amistre64):

sec = hyp/adj

OpenStudy (amistre64):

its cos upside down

OpenStudy (anonymous):

ya thats right

OpenStudy (amistre64):

\[1/\sqrt{4t^2 +1}\]

OpenStudy (anonymous):

if its hyp/adj shouldnt it just be \[\sqrt{4t^2+1}\]

OpenStudy (amistre64):

.... i spose if you want to be absolutley technical lol

OpenStudy (amistre64):

[ln(sqrt(17) + 4)] - [ln(1)]

OpenStudy (amistre64):

ummm....4(4) = 16 + 1 = 17 right?

OpenStudy (anonymous):

ya

OpenStudy (amistre64):

and the other side is sqrt(4(0) + 1) + 2(0) = ln(1)

OpenStudy (amistre64):

i hope you got an answer key, cause id love to know if were right on this :)

OpenStudy (amistre64):

2.09 is what I get....